OpenStudy (anonymous):

Find the slope of the line containing the given pair of points. If the slope is undefined state this. (-2,10) and (5,4)

OpenStudy (anonymous):

m=-6/7

OpenStudy (anonymous):

y2-y1/x2-x1 take 4-10=-6 then 5+2=7

OpenStudy (anonymous):

the slope is 7?

OpenStudy (anonymous):

 this is the slope

OpenStudy (anonymous):

 this is the formula

OpenStudy (anonymous):

ok